# Character Complexity: A Novel Measure for Quantum Circuit Analysis

### v1

@@ 1 Introduction
    Theoretical Bounds: Character Complexity ($C(U)$) is bounded by $0\leq C(U)\leq\min(n,\log_{2}(\operatorname{rank}(U)))$, where $n$ is the number of qubits.
@@ 3 Character Complexity: Definition and Properties
Let $U$ be a unitary operation representing an $n$-qubit quantum circuit, and let $G$ be the finite group generated by the gates in $U$. The character complexity $C(U)$ is bounded as follows:
$0\leq C(U)\leq\min(n,\log_{2}(\operatorname{rank}(U)))$
where $\operatorname{rank}(U)$ is the rank of the matrix representation of $U$.
@@ 1. Lower bound: $0\leq C(U)$
2. Upper bound: $C(U)\leq\min(n,\log_{2}(\operatorname{rank}(U)))$
We prove this in two parts:
a) $C(U)\leq n$:
Using the fact that $\sum_{i}d_{i}^{2}=|G|$ [10], we apply the Cauchy-Schwarz inequality:
$\sum_{i}\frac{|\chi_{i}(U)|^{2}}{d_{i}}$ $\leq\sqrt{\sum_{i}|\chi_{i}(U)|^{2}}\sqrt{\sum_{i}\frac{1}{d_{i}^{2}}}$
$\leq\sqrt{|G|}\sqrt{\sum_{i}\frac{1}{d_{i}^{2}}}\leq\sqrt{|G|}\sqrt{\sum_{i}1}=\sqrt{|G|}\sqrt{2^{n}}$
The last inequality uses the fact that $d_{i}\geq 1$ for all $i$, and there are at most $2^{n}$ irreducible representations for an $n$-qubit system.
Therefore,
$C(U)=\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(U)|^{2}}{d_{i}}\right)\leq\frac{\sqrt{|G|}\sqrt{2^{n}}}{|G|}=\sqrt{\frac{2^{n}}{|G|}}\leq n$
The last inequality holds because $|G|\geq 2^{n}$ for any group representing $n$-qubit operations, as the group must be able to generate at least $2^{n}$ distinct states.
b) $C(U)\leq\log_{2}(\operatorname{rank}(U))$:
We use the following results from representation theory:
1) The character orthogonality relation [11]:
@@ 2. Upper bound: $C(U)\leq\min(n,\log_{2}(\operatorname{rank}(U)))$
2) A corollary of Schur’s lemma [12]:
|$|\{g\in G:gU=Ug\}|\leq\operatorname{rank}(U)$
@@
$\sum_{i}\frac{|\chi_{i}(U)|^{2}}{d_{i}}$ $\leq\sum_{i}|\chi_{i}(U)|^{2}=|G|\cdot|\{g\in G:gU=Ug\}|\leq|G|\cdot\operatorname{rank}(U)$
Note that we can drop the $1/d_{i}$ term in the first inequality because $\chi_{i}(U)$ is a character of a unitary matrix, which implies $|\chi_{i}(U)|\leq d_{i}$.
Therefore,
$C(U)=\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(U)|^{2}}{d_{i}}\right)\leq\operatorname{rank}(U)$
Taking the logarithm of both sides (which preserves the inequality as $\log_{2}$ is monotonically increasing):
$C(U)\leq\log_{2}(\operatorname{rank}(U))$
Combining the results from (a) and (b), we conclude:
$C(U)\leq\min(n,\log_{2}(\operatorname{rank}(U)))$
@@
Intuitively, the character complexity $C(U)$ measures how ”spread out” the action of $U$ is across different irreducible representations of the group $G$. A low value of $C(U)$ suggests that $U$’s action is concentrated in a few representations, indicating a simpler or more structured operation. Conversely, a high value of $C(U)$ (approaching the upper bound) indicates that $U$’s action is distributed across many representations, suggesting a more complex or less structured operation. This measure provides a way to quantify the complexity of quantum operations that goes beyond simple gate counts or circuit depths, offering insights into the underlying structure of quantum algorithms.
@@ 6 Comparison of Different Quantum Circuits
  • Random Circuit (Figure 2(a)): The distribution for the random circuit shows a broad range of complexity values with a noticeable peak in the middle range. This suggests that random circuits explore a variety of complexity states, with a tendency towards moderate complexity values.
  • Quantum Fourier Transform (Figure 2(b)): The QFT displays a more structured distribution with a clear peak at moderate complexity values and a smaller peak at higher complexities. This bimodal distribution could be related to the periodic nature of the Fourier transform and suggests that QFT operates within specific complexity regimes.
  • Quantum Approximate Optimization Algorithm (Figure 2(c)): QAOA shows a unique distribution with a pronounced peak at moderate complexity values and a tail extending towards higher complexity. This distribution might reflect QAOA’s iterative approach to optimization, where the algorithm explores a specific subspace of the Hilbert space relevant to the problem being solved.
@@
  • Random circuits (Figure 3(a)) show distinct bands of increasing complexity with qubit count. The dependence on gate count is less pronounced but still visible, particularly for lower qubit counts.
  • QFT (Figure 3(b)) exhibits clear bands of increasing complexity with qubit count. The dependence on gate count is less pronounced, suggesting that the QFT’s complexity is more strongly influenced by the number of qubits than the number of gates.
  • QAOA (Figure 3(c)) shows a strong dependence on qubit count, with distinct bands of increasing complexity. The dependence on gate count is less pronounced but still visible, particularly for higher qubit counts.
@@
  • Random circuits (Figure 4(a)) display a rapid initial increase in complexity followed by saturation. This suggests that beyond a certain point, adding more gates to a random circuit does not significantly increase its complexity.
  • QFT (Figure 4(b)) shows a step-like increase in complexity. This could reflect the structured nature of the QFT algorithm, where complexity increases occur at specific points in the circuit construction.
  • QAOA (Figure 4(c)) exhibits a similar step-like increase to QFT, but with more pronounced steps. This could correspond to the iterative structure of the QAOA algorithm, where each iteration potentially introduces a significant increase in complexity.
These comparisons highlight how character complexity can reveal fundamental differences and similarities in the structure and behavior of various quantum algorithms. The unique scaling behaviors and complexity distributions observed for different circuit types underscore the richness and diversity of the quantum algorithmic landscape, and hint at the potential for character complexity to help researchers navigate this landscape more effectively.
@@
    Algorithm Design and Optimization: The distinct complexity profiles of different algorithms could guide the design of more efficient quantum circuits and inform optimization strategies.
@@
    Quantum-Classical Boundaries: The complexity scaling behaviors could provide insights into the transition between classically simulable and quantum-advantageous regimes for different algorithms.
@@
    Benchmarking: Character complexity could serve as a new metric for benchmarking quantum algorithms and hardware, complementing existing measures like circuit depth and gate count.
@@
    Error Correction Strategies: The complexity landscapes revealed by these visualizations could inform the development of error correction strategies tailored to specific algorithm types or complexity regimes.
By providing a new lens through which to view quantum algorithms, character complexity and its visualizations open up exciting avenues for future research in quantum computing theory and practice.

### v2

@@ 1 Introduction
    Theoretical Bounds: Character Complexity ($C(U)$) is bounded by $0\leq C(U)\leq\min(n,\operatorname{rank}(U))$, where $n$ is the number of qubits.
@@ 3 Character Complexity: Definition and Properties
Let $U$ be a unitary operation and $G$ be a finite group. The character complexity $C(U)$ is bounded as follows:
$0\leq C(U)\leq\sqrt{\frac{k}{|G|}}\leq 1$
where $k$ is the number of irreducible representations of $G$.
@@ 1. Lower bound: $0\leq C(U)$
2. Upper bound: $C(U)\leq\sqrt{\frac{k}{|G|}}\leq 1$
We start by applying the Cauchy-Schwarz inequality:
$\sum_{i}\frac{|\chi_{i}(U)|^{2}}{d_{i}}$ $\leq\sqrt{\sum_{i}|\chi_{i}(U)|^{2}}\sqrt{\sum_{i}\frac{1}{d_{i}}}$
We use the character orthogonality relation [11]:
@@
Next, we apply the Cauchy-Schwarz inequality to the dimension sum formula $\sum_{i}d_{i}^{2}=|G|$ [10]:
$\sum_{i}\frac{1}{d_{i}}\leq\sqrt{\sum_{i}1}\sqrt{\sum_{i}\frac{1}{d_{i}^{2}}}=\sqrt{k}\sqrt{\frac{1}{|G|}}$
@@
$C(U)$ $=\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(U)|^{2}}{d_{i}}\right)$
$\leq\frac{1}{|G|}\sqrt{|G|\cdot|\{g\in G:gU=Ug\}|}\sqrt{k}\sqrt{\frac{1}{|G|}}$
=|$=\sqrt{\frac{|\{g\in G:gU=Ug\}|}{|G|}}\sqrt{\frac{k}{|G|}}$
Since |$|\{g\in G:gU=Ug\}|$ is the order of a subgroup of $G$, we have |$|\{g\in G:gU=Ug\}|\leq|G|$. Therefore:
$C(U)\leq\sqrt{\frac{k}{|G|}}$
Finally, since $k\leq|G|$ for any finite group (as the number of irreducible representations cannot exceed the order of the group), we have:
$\sqrt{\frac{k}{|G|}}\leq\sqrt{\frac{|G|}{|G|}}=1$
@@
The character complexity $C(U)$ measures how ”spread out” the action of $U$ is across different irreducible representations of the group $G$. A low value of $C(U)$ suggests that $U$’s action is concentrated in a few representations, indicating a simpler or more structured operation. Conversely, a value of $C(U)$ approaching $\sqrt{\frac{k}{|G|}}$ indicates that $U$’s action is more evenly distributed across many representations, suggesting a more complex or less structured operation.
Values will generally fall between 0 and $\sqrt{\frac{k}{|G|}}$, with values close to $\sqrt{\frac{k}{|G|}}$ indicating operations that maximally spread their action across all irreducible representations, and values close to 0 suggesting operations that act more trivially. This measure provides a way to quantify the complexity of operations that goes beyond simple counts, offering insights into the underlying structure of algorithms by emphasizing the relative distribution of an operation’s action across the irreducible representations.
@@ 6 Comparison of Different Quantum Circuits
  • Random Circuit (Figure 2(a)): The distribution for the random circuit shows a concentration of complexity values towards the lower end of the spectrum. This suggests that random circuits tend to introduce significant scrambling, leading to lower character complexity values.
  • Quantum Fourier Transform (Figure 2(b)): The QFT displays a more structured distribution with a clear peak at moderate complexity values. This indicates that QFT, while introducing some scrambling, preserves more of the circuit’s inherent structure compared to random circuits.
  • Quantum Approximate Optimization Algorithm (Figure 2(c)): QAOA shows a distribution shifted towards higher complexity values compared to random circuits. This might reflect QAOA’s iterative approach to optimization, where the algorithm gradually introduces scrambling while maintaining some structure related to the problem being solved.
@@
  • Random circuits (Figure 3(a)) show a general trend of decreasing complexity with increasing qubit and gate counts. This aligns with the idea that more qubits and gates provide more opportunities for scrambling.
  • QFT (Figure 3(b)) exhibits a similar trend to random circuits, but with generally higher complexity values. This suggests that QFT, while introducing scrambling, maintains more structure than purely random circuits.
  • QAOA (Figure 3(c)) shows a less pronounced decrease in complexity with increasing qubit and gate counts compared to random circuits. This could reflect QAOA’s balance between scrambling and preserving problem-related structure.
@@
  • Random circuits (Figure 4(a)) display a rapid decrease in complexity with increasing gate count. This suggests that adding more gates to a random circuit quickly introduces significant scrambling.
  • QFT (Figure 4(b)) shows a more gradual decrease in complexity compared to random circuits. This could reflect the structured nature of the QFT algorithm, where scrambling is introduced more slowly as the circuit grows.
  • QAOA (Figure 4(c)) exhibits a slower decrease in complexity compared to both random circuits and QFT. This could correspond to QAOA’s iterative approach, where each iteration introduces some scrambling while maintaining problem-related structure.
These comparisons highlight how character complexity can reveal fundamental differences and similarities in the scrambling behavior of various quantum algorithms. The unique scaling behaviors and complexity distributions observed for different circuit types underscore the diversity of scrambling dynamics in the quantum algorithmic landscape.
@@
    Scrambling and Quantum Advantage: The relationship between scrambling (as measured by character complexity) and quantum advantage could be investigated. Do algorithms that exhibit more scrambling offer greater potential for speedup over classical methods?
@@
    Optimal Scrambling: The character complexity profiles could guide the design of algorithms that achieve an optimal balance between scrambling (for computational power) and structure preservation (for specific problem solving).
@@
    Scrambling and Error Correction: The interplay between scrambling and error correction could be explored. How does the introduction of error correction affect the scrambling behavior of different algorithms?
@@
    Scrambling in Noisy Intermediate-Scale Quantum (NISQ) Algorithms: Character complexity could be used to study the scrambling behavior of NISQ algorithms, which operate under the constraints of current quantum hardware. This could provide insights into the potential and limitations of these algorithms.
By interpreting character complexity as a measure of scrambling, we gain a new perspective on the behavior and properties of quantum algorithms. This opens up exciting avenues for future research at the intersection of quantum complexity, quantum chaos, and quantum algorithm design.